OpenStudy (anonymous):

Simplify radical -8 + radical -6

OpenStudy (anonymous):

\[\sqrt{-8} + \sqrt{-6}\]

OpenStudy (anonymous):

Hint:\[\cdots=i\sqrt 8 + i \sqrt 6\]